# Large Language Models are In-Context Molecule Learners

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in biochemical tasks, especially the molecule caption translation task, which aims to bridge the gap between molecules and natural language texts. However, previous methods in adapting LLMs to the molecule-caption translation task required extra domain-specific pre-training stages, suffered weak alignment between molecular and textual spaces, or imposed stringent demands on the scale of LLMs. To resolve the challenges, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm allowing LLMs to learn the molecule-text alignment from context examples via In-Context Molecule Tuning. Specifically, ICMA incorporates the following three stages: Cross-modal Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Initially, Cross-modal Retrieval utilizes BM25 Caption Retrieval and Molecule Graph Retrieval to retrieve informative context examples. Additionally, we also propose Post-retrieval Re-ranking with Sequence Reversal and Random Walk to further improve the quality of retrieval results. Finally, In-Context Molecule Tuning unlocks the in-context molecule learning capability of LLMs with retrieved examples and adapts the parameters of LLMs for the molecule-caption translation task. Experimental results demonstrate that ICMT can empower LLMs to achieve state-of-the-art or comparable performance without extra training corpora and intricate structures, showing that LLMs are inherently in-context molecule learners.
@@ 1 Introduction
Molecules play a crucial role across various fields, such as medicine  Ding et al. (2019), agriculture  Twyman et al. (2003), and material science  Higuchi et al. (2023), as they are widely used in the development of drugs, fertilizers, and advanced materials. Recently, LLMs have demonstrated remarkable success in the molecular domain, as molecules can be represented as Simplified Molecular-Input Line-Entry System (SMILES) strings  Weininger (1988), which can be comprehended and generated by LLMs in a similar manner to natural languages. To further bridge the gap between the molecules and natural languages,  Edwards et al. (2022) proposes the molecule-caption translation task, which comprises two sub-tasks: molecule captioning (Mol2Cap) and text-based de novo molecule generation (Cap2Mol). Specifically, Mol2Cap involves generating a textual description that elucidates the features of the given molecule, while Cap2Mol focuses on predicting the exact molecule based on the textual caption. The study of the molecule-caption translation task offers an accessible and user-friendly venue for molecule discovery, which has raised wide research focus.
@@
Generally, there are two main paradigms for adapting LLMs to the molecule-caption retrieval task. The first paradigm is the domain-specific pre-training & fine-tuning. For instance, MolT5  Edwards et al. (2022) first proposes and handles the molecule-caption translation task as the language translation task, pre-training the MolT5 model with chemical corpora like PubChem  Kim et al. (2019) and then fine-tuning the model on the ChEBI-20 dataset  Edwards et al. (2022). Additionally, MoMu  Su et al. (2022) and MolCA  Liu et al. (2023b) introduce an extra pre-training stage before fine-tuning on downstream tasks, which aligns the output of 2D molecule graph encoder with the input space of LLMs. In contrast, the other paradigm involves prompting and utilizing the in-context learning capability of LLMs. For example, MolReGPT  Li et al. (2023) introduces In-Context Few-Shot Molecule Learning, prompting general LLMs like GPT-3.5 and GPT-4 to achieve competitive performance without extra parameter adaptations.
@@
However, the current paradigms still face challenges. On one hand, the domain-specific pre-training & fine-tuning paradigm requires extra pre-training stages, which is challenging due to the scarcity of high-quality chemical datasets, especially molecule-caption pairs, making this paradigm difficult to scale up to the most advanced LLMs with billion parameters. Besides, the domain-specific pre-training & fine-tuning paradigm also suffers from weak alignment between molecules and texts, as phrases in molecule captions often describe specific sub-structures of molecules rather than the entire molecule. Despite attempts trying to introduce extra modalities for better alignment, the integration of the additional modalities (e.g., 2D molecule graph) is still focused on the entire graph-level and can only be applied to the molecule captioning task, while ignoring the text-based generation of molecules, which is much more valuable for drug discovery. On the other hand, the in-context learning & prompting paradigm puts a harsh requirement on LLMs’ reasoning and in-context learning capabilities, which tends to emerge when LLMs are large enough, making them computationally infeasible. Consequently, there is a demand for a unified and efficient approach that effectively enhances the performance of the most advanced LLMs in both two sub-tasks of molecule-caption translation.
@@
In this case, we propose In-Context Molecule Adaptation (ICMA) as a new paradigm for adapting LLMs in molecule-caption translation. Different from previous paradigms, ICMA aims to instruct LLMs to derive knowledge from informative context examples, especially the alignment between molecule SMILES representations and captions, via In-Context Molecule Tuning. As shown in Figure 1, similar molecules often share similar properties, as indicated by the overlaps among molecule captions. Conversely, similar captions tend to describe molecules with similar SMILES representations. In this case, with ICMA, general LLMs could fulfil their reasoning and in-context learning capability to better grasp the alignment between molecules and textual captions from context examples, thereby achieving better performance.
@@
Specifically, ICMA incorporates three stages: Cross-modal Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. In the initial stage, Cross-modal Retrieval, we employ Caption Retrieval and Molecule Graph Retrieval to fetch similar captions and molecules, respectively. Subsequently, we introduce the Post-retrieval Re-ranking stage to enhance the quality of the retrieval algorithms. This stage incorporates two innovative strategies: Sequence Reversal and Random Walk, which aim to refine and prioritize the retrieved examples. Finally, we apply In-context Molecule Tuning to adapt the parameters of LLMs, enabling them to learn from the contextual mappings and effectively utilize the learned knowledge for the current generation task. Experiments are conducted across two real-world molecule-caption translation datasets, ChEBI-20 and PubChem324k. Results show that ICMA could enable LLMs to achieve state-of-the-art or comparable performance in both the two sub-tasks (i.e., Mol2Cap and Cap2Mol). Meanwhile, we also study the factors related to the model performance, including retrieval algorithms, context settings (i.e., context example number and maximum input length), and model scales. Lastly, ablation studies are conducted to justify the effectiveness of Post-retrieval Re-ranking components.
@@
  • We implement ICMA through three stages, including Cross-modal Retrieval, Post-retrieval Re-ranking, and In-Context Molecule Tuning, significantly enhancing the informativeness of context examples.
@@ 2.1 Molecule-Caption Translation
The molecule-caption translation task is initially proposed in MolT5 Edwards et al. (2022). Inspired by the image captioning task, MolT5 introduces a new dataset, ChEBI-20, with pairs of molecules and manually labelled captions that describe the molecular properties. Meanwhile, MolT5 proposes a T5 model that is jointly pre-trained on molecule SMILES and general text corpus. MolXPT  Liu et al. (2023a) pre-trains a GPT model by introducing extra-wrapped texts as the pre-training corpus, demonstrating better molecule-text alignment. However, the generation of SMILES strings suffers from the problem of invalid SMILES due to the mismatches of brackets. To overcome the generation issue of SMILES strings, BioT5  Pei et al. (2023) introduces Self-referencing Embedded Strings (SELFIES)  Krenn et al. (2020) instead of SMILES strings to represent molecules in LLMs and proposes the BioT5 model that is jointly trained on single-modal data, wrapped text data, and molecule/protein-description pairs. Meanwhile, as molecules can also be represented as graphs, some methods focus on molecule understanding by introducing molecule graph information. For example, MoMu  Su et al. (2022) first proposes a graph encoder to encode molecule graph information and utilizes contrastive learning to bridge the semantic gap between the graph encoder and the LLM. To better fuse the molecule graph information, MolCA  Liu et al. (2023b) follows the BLIP-2 and utilizes a Q-Former to project the output of the graph encoder into the LLMs, showing better molecule understanding performance. However, most of these methods still adhere to the pre-training & fine-tuning paradigm, which necessitates the joint pre-training of LLMs on both general text and extra chemical domain corpora. As the size of the training corpora and model weights of LLMs continue to increase, this approach has become extremely inefficient. To address this issue, MolReGPT  Li et al. (2023) proposes the In-Context Few-Shot Molecule Learning to enable LLMs to learn the molecule-caption translation task from the context examples without modifying the model weights while still achieving comparable performance to these fine-tuned methods.
@@ 2.2 In-Context Learning
With the scaling of model size and corpus size  Brown et al. (2020); Chowdhery et al. (2023), LLMs emerge the in-context learning capability  Wang et al. (2022), which enables LLMs to learn from contexts augmented with several examples  Dong et al. (2022). By utilizing the capability of ICL, LLMs can solve complex tasks without the necessity of being fine-tuned. For instance, with a few examples, GPT-3 could demonstrate similar performance to fine-tuned models in unseen tasks  Brown et al. (2020). What’s more, based on context examples, LLMs could achieve better mathematical reasoning ability with the assistance of chain-of-thought (CoT)  Wei et al. (2022). In the scenario of molecule-caption translation, MolReGPT  Li et al. (2023) proves the retrieval quality is closely related to the model performance, while it still imposes strict requirements on the model scale, as the in-context learning capability only becomes apparent when the model weights reach a certain size.
@@ 3 In-Context Molecule Adaptation
Figure 2: Framework of In-Context Molecule Adaptation (ICMA). Generally, ICMA consists of three stages, Cross-modal Retrieval, Post-retrieval Re-ranking, and In-Context Molecule Tuning.
In this section, we introduce In-Context Molecule Adaptation (ICMA) as a novel paradigm to adapt LLMs to molecule-caption translation. As shown in Figure 2, ICMA incorporates three stages, including Cross-modal Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Specifically, Cross-modal Retrieval first retrieves $N$ rough examples from the training set $\mathcal{D}$ by calculating the similarity between the current query and molecule-caption pairs. After that, Post-retrieval Re-ranking with Random Walk and Sequence Reversal is adopted to obtain $n$ refined examples from the $N$ rough examples. Finally, the In-context Molecule Tuning can be performed to update the parameters of LLMs to learn the molecule-text alignment from refined context examples.
3.1 Cross-modal Retrieval
The retrieval quality is closely related to the informativeness of context examples. For example, If the retrieved molecules are more similar to the current query molecule, they are likely to exhibit more overlaps in their respective caption, which could enable better alignments between molecules and texts. Therefore, the development of retrieval algorithms plays a crucial role in ICMA.
For the Mol2Cap task, ICMA adopts Molecule Graph Retrieval to better refine the retrieval quality. Previously, MolReGPT  Li et al. (2023) utilizes the Morgan Fingerprints (Morgan FTS) and Dice similarity to evaluate the similarity between molecules, which encodes the pre-defined handcraft structures as the embedding of the molecule. However, Morgan FTS are typically based on handcrafted chemical feature extraction, which may have limited capability to capture the comprehensive information of complex molecule structures and properties. On the other side, Graph Neural Networks (GNNs), pre-trained on millions of molecules, could effectively capture topological structures of molecule graphs, providing complete chemical semantics Xia et al. (2023). This makes GNNs a better option for molecule similarity calculation. In ICMA, we adopt a pre-trained GNN encoder to obtain the molecule graph embeddings
@@ 3.1 Cross-modal Retrieval
In the Cap2Mol task, we inherit the BM25 Caption Retrieval from MolReGPT  Li et al. (2023) as it focuses on the detail matching of molecule captions, showing competitive performance and is much faster than LLM-based methods like Sentencebert  Reimers and Gurevych (2019). Specifically, given captions in the test set as query captions $Q_{c}$ and the training set of captions $\mathcal{D}_{c}$, the caption similarity ranking function $\mathcal{R}^{c}$ can be denoted as:
@@ 3.2.1 Random Walk
The molecules ranked top by retrieval algorithms can sometimes share too many overlaps, impairing the information density. In this case, for the context diversity and generalization performance of ICMA, it is necessary to give these less similar examples a chance to be visited. Inspired by the random walk mechanism in graph theory, we propose random walk as a post-retrieval method to select examples from the top-$N$ retrieved results so that examples with lower rank still have a chance to be selected, allowing the complement of retrieval algorithms Mathematically, we adopt a dynamic chance for examples with different ranks. Specifically, for the $j$-th example in the $N$ rouge results, where $1\leq j\leq$N, the possibility of skipping is represented as $(10-j)\%$, which reduces as the rank moves down to ensure that at least one example is selected for the context. If the $j$-th example is skipped, then it will perform Random Walk in the remaining examples until an example is selected.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in biochemical tasks, especially the molecule caption translation task, which aims to bridge the gap between molecules and natural language texts. However, previous methods in adapting LLMs to the molecule-caption translation task required extra domain-specific pre-training stages, suffered weak alignment between molecular and textual spaces, or imposed stringent demands on the scale of LLMs. To resolve the challenges, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm allowing LLMs to learn the molecule-text alignment from context examples via In-Context Molecule Tuning. Specifically, ICMA incorporates the following three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Initially, Hybrid Context Retrieval utilizes BM25 Caption Retrieval and Molecule Graph Retrieval to retrieve informative context examples. Additionally, we also propose Post-retrieval Re-ranking with Sequence Reversal and Random Walk to further improve the quality of retrieval results. Finally, In-Context Molecule Tuning unlocks the in-context molecule learning capability of LLMs with retrieved examples and adapts the parameters of LLMs for the molecule-caption translation task. Experimental results demonstrate that ICMA can empower LLMs to achieve state-of-the-art or comparable performance without extra training corpora and intricate structures, showing that LLMs are inherently in-context molecule learners.
@@ 1 Introduction
Molecules play a crucial role across various fields, such as medicine  Ding et al. (2019), agriculture  Twyman et al. (2003), and material science  Higuchi et al. (2023), as they are widely used in the development of drugs, fertilizers, and advanced materials. Recently, LLMs have demonstrated remarkable success in the molecular domain, as molecules can be represented as Simplified Molecular-Input Line-Entry System (SMILES) strings  Weininger (1988), which can be comprehended and generated by LLMs in a similar manner to natural languages. To further bridge the gap between the molecules and natural languages,  Edwards et al. (2022) proposes the molecule-caption translation task, which comprises two sub-tasks: molecule captioning (Mol2Cap) and text-based de novo molecule generation (Cap2Mol). Specifically, Mol2Cap involves generating a textual description that elucidates the features of the given molecule, while Cap2Mol focuses on predicting the exact molecule based on the textual caption. The study of the molecule-caption translation task offers an accessible and chemist-friendly venue for molecule discovery, which has raised wide research focus.
@@
Generally, there are two main paradigms for adapting LLMs to the molecule-caption retrieval task. The first paradigm is the domain-specific pre-training & fine-tuning. For instance, MolT5  Edwards et al. (2022) first proposes and handles the molecule-caption translation task as the language translation task, pre-training the MolT5 model with chemical corpora like PubChem  Kim et al. (2019) and then fine-tuning the model on the ChEBI-20 dataset  Edwards et al. (2021). Additionally, MoMu  Su et al. (2022) and MolCA  Liu et al. (2023b) introduce an extra modality alignment stage before fine-tuning on downstream tasks, which aligns the output of 2D molecule graph encoder with the input space of LLMs. In contrast, the other paradigm involves prompting and utilizing the in-context learning capability of LLMs. For example, MolReGPT  Li et al. (2023) introduces In-Context Few-Shot Molecule Learning, prompting general LLMs like GPT-3.5 and GPT-4 to achieve competitive performance without extra parameter adaptations.
@@
However, the current paradigms still face challenges. On one hand, the domain-specific pre-training & fine-tuning paradigm requires extra pre-training stages (i.e., domain-specific pre-training and modality alignment), which is challenging due to the scarcity of high-quality chemical datasets, especially molecule-caption pairs, making this paradigm difficult to scale up to the most advanced LLMs with billion parameters. Besides, the domain-specific pre-training & fine-tuning paradigm also suffers from weak alignment between molecules and texts, as phrases in molecule captions often indicate specific sub-structures of molecules rather than the entire molecule. Despite attempts to introduce extra modalities for better alignment Su et al. (2022); Liu et al. (2023b), the integration of the additional modalities (e.g., 2D molecule graph) is still focused on the entire graph level and can only be applied to the molecule captioning task, while ignoring the text-based generation of molecules, which is much more valuable for drug discovery. On the other hand, the in-context learning & prompting paradigm puts a harsh requirement on LLMs’ emergent capabilities, such as reasoning and in-context learning abilities. However, LLMs with these emergent capabilities usually have billions of parameters, making them computationally expensive. Consequently, there is a demand for a unified and efficient approach that effectively enhances the performance of the most advanced LLMs in both two sub-tasks of molecule-caption translation.
@@
In this case, we propose In-Context Molecule Adaptation (ICMA) as a new paradigm for adapting LLMs in molecule-caption translation. Different from previous paradigms, ICMA aims to instruct LLMs to derive knowledge from informative context examples, especially the alignment between molecule SMILES representations and captions, via In-Context Molecule Tuning. As shown in Figure 1, similar molecules often share similar properties, as indicated by the overlaps among molecule captions. Conversely, similar captions tend to describe molecules with similar SMILES representations. In this case, with ICMA, general LLMs could fulfill their reasoning and in-context learning capability to better grasp the alignment between molecules and textual captions from context examples, thereby achieving better performance.
@@
Specifically, ICMA incorporates three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. In the initial stage, Hybrid Context Retrieval, we employ Caption Retrieval and Molecule Graph Retrieval to fetch similar captions and molecules, respectively. Subsequently, we introduce the Post-retrieval Re-ranking stage to enhance the quality of the retrieval algorithms. This stage incorporates two innovative strategies: Sequence Reversal and Random Walk, which aim to refine and reprioritize the retrieved examples. Finally, we apply In-context Molecule Tuning to adapt the parameters of LLMs, enabling them to learn from the contextual mappings and effectively utilize the learned knowledge for the current generation task. Experiments are conducted across two real-world molecule-caption translation datasets, ChEBI-20 and PubChem324k. Results show that ICMA could enable LLMs to achieve state-of-the-art or comparable performance in both the two sub-tasks (i.e., Mol2Cap and Cap2Mol). Meanwhile, we also study the factors related to the model performance, including retrieval algorithms, context settings (i.e., context example number and maximum input length), and model scales. Lastly, ablation studies are conducted to justify the effectiveness of Post-retrieval Re-ranking components.
@@
  • We implement ICMA through three stages, including Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-Context Molecule Tuning, significantly enhancing the informativeness of context examples.
@@ 2.1 Molecule-Caption Translation
Inspired by the image captioning task, Edwards et al. (2021) introduces a new dataset, ChEBI-20, with pairs of molecules and manually labeled captions that describe the molecular properties. The molecule-caption translation task was initially proposed in MolT5 Edwards et al. (2022). Meanwhile, MolT5 proposes a T5 model that is jointly pre-trained on molecule SMILES and general text corpus. MolXPT  Liu et al. (2023a) pre-trains a GPT model by introducing extra-wrapped texts as the pre-training corpus, demonstrating better molecule-text alignment. However, the generation of SMILES strings suffers from the problem of invalid SMILES due to the mismatches of brackets. To overcome the generation issue of SMILES strings, BioT5  Pei et al. (2023) introduces Self-referencing Embedded Strings (SELFIES)  Krenn et al. (2020) instead of SMILES strings to represent molecules in LLMs and proposes the BioT5 model that is jointly trained on single-modal data, wrapped text data, and molecule/protein-description pairs. Meanwhile, as molecules can also be represented as graphs, some methods focus on molecule understanding by introducing molecule graph information. For example, MoMu  Su et al. (2022) first proposes a graph encoder to encode molecule graph information and utilizes contrastive learning to bridge the semantic gap between the graph encoder and the LLM. To better fuse the molecule graph information, MolCA  Liu et al. (2023b) follows the BLIP-2 and utilizes a Q-Former to project the output of the graph encoder into the LLMs, showing better molecule understanding performance. However, most of these methods still adhere to the pre-training & fine-tuning paradigm, which necessitates the joint pre-training of LLMs on both general text and extra chemical domain corpora. As the size of the training corpora and model weights of LLMs continue to increase, this approach has become extremely inefficient. To address this issue, MolReGPT  Li et al. (2023) proposes the In-Context Few-Shot Molecule Learning to enable LLMs to learn the molecule-caption translation task from the context examples without modifying the model weights while still achieving comparable performance to these fine-tuned methods.
@@ 2.2 In-Context Learning
With the scaling of model size and corpus size  Brown et al. (2020); Chowdhery et al. (2023), LLMs emerge the in-context learning capability  Wang et al. (2022), which enables LLMs to learn from contexts augmented with several examples  Dong et al. (2022). By utilizing the capability of ICL, LLMs can solve complex tasks without the necessity of being fine-tuned. For instance, with a few examples, GPT-3 could demonstrate similar performance to fine-tuned models in unseen tasks  Brown et al. (2020). What’s more, based on context examples, LLMs could achieve better mathematical reasoning ability with the assistance of chain-of-thought (CoT)  Wei et al. (2022). With the powerful in-context learning capabilities,  Edwards et al. (2023) proposes in-context drug synergy learning to apply LLMs for personalized drug synergy prediction and drug design, while  Jablonka et al. (2024) fine-tunes LLMs such as GPT-3 with chemical questions and answers to solve predictive tasks In the scenario of molecule-caption translation, MolReGPT  Li et al. (2023) proves the retrieval quality is closely related to the model performance, while it still imposes strict requirements on the model scale, as the in-context learning capability only becomes apparent when the model weights reach a certain size.
@@ 3 In-Context Molecule Adaptation
Figure 2: Framework of In-Context Molecule Adaptation (ICMA). Generally, ICMA consists of three stages, Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-Context Molecule Tuning.
In this section, we introduce In-Context Molecule Adaptation (ICMA) as a novel paradigm to adapt LLMs to molecule-caption translation. As shown in Figure 2, ICMA incorporates three stages, including Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Specifically, Hybrid Context Retrieval first retrieves $N$ rough examples from the training set $\mathcal{D}$ by calculating the similarity between the current query and molecule-caption pairs. After that, Post-retrieval Re-ranking with Random Walk and Sequence Reversal is adopted to obtain $n$ refined examples from the $N$ rough examples. Finally, the In-context Molecule Tuning can be performed to update the parameters of LLMs to learn the molecule-text alignment from refined context examples.
3.1 Hybrid Context Retrieval
The retrieval quality is closely related to the informativeness of context examples. For example, if the retrieved molecules are more similar to the current query molecule, they are likely to exhibit more overlaps in their respective caption, which could enable better alignments between molecules and texts. Therefore, the development of retrieval algorithms plays a crucial role in ICMA. In this work, we introduce Hybrid Context Retrieval, which adopts hybrid modalities (i.e., 2D molecule graph and text), as well as hybrid retrieval algorithms designed for the specific tasks (i.e., Molecule Graph Retrieval for Mol2Cap and BM25 Caption Retrieval for Cap2Mol).
For the Mol2Cap task, ICMA adopts Molecule Graph Retrieval to better refine the retrieval quality. Previously, MolReGPT  Li et al. (2023) utilizes the Morgan Fingerprints (Morgan FTS) and Dice similarity to evaluate the similarity between molecules, which encodes the pre-defined handcraft structures as the embedding of the molecule. However, Morgan FTS are typically based on handcrafted chemical feature extraction, which may have limited capability to capture the comprehensive information of complex molecule structures and properties. On the other side, Graph Neural Networks (GNNs), pre-trained on millions of molecules, could effectively capture topological structures of molecule graphs, providing complete chemical semantics Xia et al. (2023). This makes GNNs a better option for molecule similarity calculation. In ICMA, we adopt a pre-trained GNN encoder to obtain the molecule graph embeddings:
@@ 3.1 Hybrid Context Retrieval
In the Cap2Mol task, we inherit the BM25 Caption Retrieval  Robertson et al. (2009) from MolReGPT  Li et al. (2023) as it focuses on the detail matching of molecule captions, showing competitive performance and is much faster than LLM-based methods like Sentencebert  Reimers and Gurevych (2019). Specifically, given captions in the test set as query captions $Q_{c}$ and the training set of captions $\mathcal{D}_{c}$, the caption similarity ranking function $\mathcal{R}^{c}$ can be denoted as:
@@ 3.2.1 Random Walk
The molecules ranked top by retrieval algorithms can sometimes share too many overlaps, impairing the informativeness of context examples. In this case, for the context diversity and generalization performance of ICMA, it is necessary to give these less similar examples a chance to be visited. Inspired by the random walk mechanism in graph theory, we propose random walk as a post-retrieval method to select examples from the top-$N$ retrieved results so that examples with lower rank still have a chance to be selected, which provides more useful information and complements the context diversity. Mathematically, we adopt a dynamic chance for examples with different ranks. Specifically, for the $j$-th example in the $N$ rough results, where $1\leq j\leq$N, the possibility of skipping is represented as $(9*\frac{N-j}{N-1})\%$, which reduces as the rank moves down to ensure that at least one example is selected for the context. If the $j$-th example is skipped, then it will perform a Random Walk in the remaining examples until an example is selected. Practically, if $N$ is too large (i.e., $N\gg n$), the retrieval quality will be impaired, while if $N$ is too small (i.e., $N\approx n$), it might hurt the diversity. Therefore, considering $n$ is not quite large and usually less than 5 due to the context length limitation, $N$ is set to 10 in this work.